OpenStudy (anonymous):

calculate sin, cos, and tan of O. opposite is 7, adjacent is 6, and opposite is unknown

OpenStudy (luigi0210):

@harsh08 Welcome to Openstudy :) You should start by drawing out your triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

Мы не можем вычислить некоторые тригонометрические значения без противоположное значение.

OpenStudy (anonymous):

this is it

OpenStudy (unklerhaukus):

\[\large\boxed{\sin\theta=\frac{\text{opposite side}}{\text{hypotenuse}}},\\\large \boxed{\cos\theta=\frac{\text{adjacent side}}{\text{hypotenuse}}},\\\large\boxed{\tan\theta=\frac{\text{opposite side}}{\text{adjacent side}}}\]

OpenStudy (anonymous):

I know this

OpenStudy (anonymous):

Все, что вам нужно сделать, это вычислить значение синуса.

OpenStudy (anonymous):

but how do i do it without an opposite side?

OpenStudy (unklerhaukus):

\[\theta\]